Doc. Math. J. DMV 139

Section 3

Number Theory and Arithmetic Algebraic Geometry

In case of several authors, Invited Speakers are marked with a ^{∗}.

Vladimir G. Berkovich: p-Adic Analytic Spaces . . . . . . . . . . . . . . . . . . II 141
Pierre Colmez: Représentations p-Adiques d’un Corps Local . . . . . . . II 153
W. Duke: Bounds for Arithmetic Multiplicities . . . . . . . . . . . . . . . . . . . . . . II 163
François Gramain: Quelques Résultats d’Indépendance Algébrique II 173
Loïc Merel: Points Rationnels et Séries de Dirichlet . . . . . . . . . . . . . . . . II 183

Shinichi Mochizuki: The Intrinsic Hodge Theory of p-Adic

Hyperbolic Curves . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 187

Hans Peter Schlickewei: The Subspace Theorem and Applications II 197

Takeshi Tsuji: p-Adic Hodge Theory in the Semi-Stable Reduction

Case . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 207

Shou-Wu Zhang: Small Points and Arakelov Theory . . . . . . . . . . . . . . . II 217

Documenta Mathematica · Extra Volume ICM 1998 · II ·

140

Documenta Mathematica · Extra Volume ICM 1998 · II ·